# RLCP: A Reinforcement Learning-based Copyright Protection Method for Text-to-Image Diffusion Model

### v2

@@ 5.2 Experiment Results
In this section, we evaluate the effectiveness of our proposed RLCP method across several dimensions. We compare our results against baseline methods, measure the impact of different proportions of copyrighted data, and analyze the trade-offs between copyright protection and image quality.
Effectiveness of copyright metric. Our first set of experiments aims to assess the reliability of our proposed copyright loss (CL) metric in distinguishing between copyrighted and non-copyrighted content.
The experimental results shown in Figure 3 were obtained using a subset of the paintings dataset(See in Figure 4), which includes works such as “The Starry Night” by Van Gogh. CLIP similarity was calculated by leveraging the CLIP model to generate semantic embeddings, followed by computing the Mean Squared Error (MSE). The L2-norm was determined by calculating the squared L2 distance between these weighted activations, while LPIPS normalized the feature dimensions of all pixels and layers to unit length, scaling each feature by a specific weight.
In Figure 3, the “query index” refers to the position of the query image in the dataset, while the “value index” represents the corresponding value or similarity score for each generated image. The heatmaps illustrate good performance by showing high similarity scores for relevant image pairs, indicating that the model successfully retrieves or generates images that closely match the queries in terms of semantic or visual content.
The resulting heatmaps demonstrate that the combined copyright loss metric, which integrates both semantic and perceptual metrics, offers a more balanced and accurate assessment of copyright infringement risks. The copyright loss heatmap, in particular, exhibits a more nuanced and effective differentiation between potentially infringing and non-infringing content, highlighting its robustness in evaluating generative model outputs.
Performance of Copyright Protection. We have two primary foundational models in our repertoire. The first one is our finetuned model based on reinforcement learning, SD-finetuned, which is employed to evaluate and assess the performance of the reinforcement learning algorithm and fine-tuned the stable diffusion model with the use of similarity metrics. This evaluation includes both the model’s ability to forget copyrighted images and generated data quality. The second foundational model is on the basic SD-XL 1.0. During the unlearning experiments conducted on these two foundational models, for each image to be forgotten, the learning rate for gradient ascent is set at 3e-4, detailed hyperparameters are listed in Table 1.
@@
The result are listed on Table 2, Table 3, Table 4. While RLCP shows relatively weaker performance on the FID, it still performs exceptionally well on the L2-norm metric. This discrepancy can be explained by the nature of FID, which measures the distribution similarity between generated and target images. Since RLCP focuses on avoiding the generation of copyrighted data, it naturally alters the distribution, leading to a higher FED score. Despite this, the L2-norm results indicate that RLCP maintains strong performance in generating non-infringing content that aligns with the original image features.
We also give examples of fine-tuned results for copyright protection on three datasets (See in Figure 5, Figure 6, Figure 7).
Impact of Copyright Data Proportion. We further explore how the proportion of copyrighted images in the training dataset affects the performance of the RLCP method. As illustrated in Figure 8, increasing the proportion of copyrighted data leads to an increase in copyright loss, while FID decreases slightly, reflecting a trade-off between protection and quality. One potential concern is whether RLCP can scale effectively when a large proportion of the training set consists of copyrighted images. The results address this concern by showing that RLCP continues to perform strongly, even with a high percentage of copyrighted images, outperforming all alternative methods.
We measured copyright loss and FID across various datasets with differing ratios of copyrighted to non-copyrighted images. As shown in Figure 8, the copyright loss also rises with more copyrighted data, yet the FID shows only a slight decrease.
While RLCP demonstrates a clear ability to reduce copyright infringement, there is an inherent trade-off between minimizing copyright loss and maintaining image quality. The results suggest that RLCP provides a favorable balance, where the reduction in copyright loss is achieved with only a minimal impact on FID. This balance is particularly evident in the results for portrait images, where RLCP achieves both low copyright loss and high FID.
6 Conclusion and Future Work
@@
In this paper, we presented a Reinforcement Learning-based Copyright Protection (RLCP) for copyright infringement in text-to-image diffusion model. RLCP proposes a copyright loss metric that mirrors legal tests used to assess substantial similarity, and then integrates this metric into a reinforcement learning framework for model fine-tuning, and the use of KL divergence to regularize and stabilize the model training process. Experiments conducted on three mixed datasets of copyright and non-copyright images show that RLCP significantly reduces the likelihood of generating infringing content while preserving the visual quality of the generated images. Our results demonstrate that balancing the proportion of copyrighted and non-copyrighted data in the training set is crucial for minimizing copyright infringement without compromising image quality. We also showed that the reward-driven RL framework effectively fine-tunes diffusion models by optimizing for both copyright compliance and data fidelity.
While our approach demonstrates promising results, there are several areas for future work: (1) Broader Application Domains: Future work could extend RLCP to other domains beyond image generation, such as text or audio generation models, where copyright concerns are equally prevalent. (2) Dynamic Dataset Management: Investigating adaptive or dynamic dataset augmentation strategies could be beneficial. As models encounter more copyright-protected data, dynamically adjusting the training process may lead to more robust copyright protection without overfitting to specific datasets.

### v3

@@ 5.2 Experiment Results
We evaluate the effectiveness of our proposed RLCP method across several dimensions. We compare our results against baselines, analysis of individual metrics, and analyze the trade-offs between copyright protection and image quality.
Effectiveness of copyright metric. To validate the reliability of our copyright loss (CL) metric, we conducted experiments on 10 copyrighted paintings. CLIP embeddings were used to compute semantic similarity via Mean Squared Error, and perceptual similarity was measured using LPIPS.
In Figure 3, each heatmap reflects the similarity scores between each pair of images, with the diagonal cells highlighting scores between identical pairs. The “query index” refers to the position of the query image in the dataset, while the “value index” represents the corresponding value or similarity score for each generated image. The first two heatmaps show the similarity scores using individual metrics, where image similarity is represented, but the clarity and emphasis on identical pairs are less pronounced. In the third heatmap, which uses the combined metric, the similarity is more distinctly captured, especially along the diagonal, where the color is noticeably deeper and more defined. This shows that the combined metric provides a clearer and more accurate representation of the similarity between the images.
| Method | Paintings | Cartoons | Portraits | |||||||||
| CLIP (%) | CL (%) | -norm (%) | FID | CLIP (%) | CL (%) | -norm (%) | FID | CLIP (%) | CL (%) | -norm (%) | FID | |
| RLCPsemantic | 47.12 | 43.78 | 55.12 | 15.0 | 47.45 | 38.12 | 47.54 | 17.6 | 47.34 | 44.78 | 53.12 | 16.8 |
| RLCPperceptual | 43.45 | 39.12 | 51.12 | 17.5 | 43.56 | 30.12 | 40.78 | 19.8 | 41.45 | 38.78 | 49.56 | 17.2 |
| RLCP | 35.12 | 28.45 | 40.12 | 18.1 | 30.34 | 20.76 | 33.21 | 20.8 | 30.23 | 25.67 | 38.34 | 17.5 |
@@
Performance of Copyright Protection. The comparison of RLCP with baselines is shown in Table 1. While RLCP shows relatively weaker performance on the FID, it still performs exceptionally well on the -norm metric. This discrepancy can be explained by the nature of FID, which measures the distribution similarity between generated and target images. Since RLCP focuses on avoiding the generation of copyrighted data, it naturally alters the distribution, leading to a higher FID score. Despite this, the -norm results indicate that RLCP maintains strong performance in generating non-infringing content that aligns with the original image features. We also give examples of fine-tuned results for copyright protection on three datasets in Figure 4.
Ablation Study. To evaluate the impact of individual metrics in the reward function, we assessed the model using only semantic similarity (CLIP) or perceptual similarity (LPIPS) into Reinforcement Learning framework, excluding the combined metric. The results are shown in Table 2.
The approach achieves higher CLIP scores compared to , indicating better semantic alignment with the target content. However, it underperforms in metrics like -norm, where Perceptual-Only outperforms Semantic-Only across all categories. This suggests that Perceptual-Only generates images with features more visually distinct from copyrighted content.
The combined metric (RLCP) balances these trade-offs, achieving the lowest CLIP and -norm scores, while also demonstrating better FID performance compared to and . Integrating both metrics provides stronger copyright protection and higher generative quality than using either metric individually.
Impact of Copyright Data Proportion. Figure 5 illustrates how varying the proportion of copyrighted data impacts RLCP’s performance. As the proportion of copyrighted data increases, copyright loss (CL) rises due to the model generating images that more closely resemble the copyrighted content. Conversely, FID decreases slightly, reflecting improved image quality.
These results address concerns about scalability, showing that RLCP performs robustly even with a high concentration of copyrighted data. A lower FID score indicates better image quality, but this comes at the cost of increased copyright loss (CL). This trade-off is particularly evident when training data contains highly similar copyrighted images.
6 Conclusion
@@
In this paper, we presented a Reinforcement Learning-based Copyright Protection (RLCP) for copyright infringement in text-to-image diffusion model. RLCP proposes a copyright loss metric that mirrors legal tests used to assess substantial similarity, and then integrates this metric into a reinforcement learning framework for model fine-tuning, and the use of KL divergence to regularize and stabilize the model training process. Experiments conducted on three mixed datasets of copyright and non-copyright images show that RLCP significantly reduces the likelihood of generating infringing content while preserving the visual quality of the generated images.